OpenStudy (anonymous):

Write the expression as a single natural logarithm. 2 ln x – 4 ln c

OpenStudy (pawanyadav):

2lnx=lnx^2 4lnc=ln c^4 And ln x^2-ln c^4=ln(x^2/c^4)

OpenStudy (pawanyadav):

By property lna - lnb=ln(a/b)

OpenStudy (pawanyadav):

Got it

OpenStudy (anonymous):

i didnt get it

OpenStudy (pawanyadav):

Do you know lna-lnb=ln(a/b)

OpenStudy (anonymous):

i got it now .. thanks man

OpenStudy (pawanyadav):

You are welcome